Q1. 20cm³ of a gaseous hydrocarbon and 150cm³ of oxygen were exploded in a closed vessel. After cooling, 110cm³ of gases remained. After absorption by concentrated NaOH solution, the volume left was 50cm³. Determine the molecular formula of the hydrocarbon
1
Expert's answer
2020-05-13T14:20:51-0400
  1. 110 cm3 - 50 cm3 = 60 cm3 it`s CO2
  2. 50 cm3 - unburnt oxygen
  3. 150 cm3 - 50 cm3 = 100 cm3 consumed oxygen
"20C_xH_y + 100 O_2 = 60 CO_2 + zH_2O"

5.

"2C_xH_y + 10O_2 = 6CO_2 + 8H_2O"

6.

"C_xH_y + 5O_2 = 3CO_2 + 4H_2O"

7.

"x = 3; y = 8."

Answer: C3H8